OpenStudy (anonymous):

jennifer earned $40 LESS THEN TWICE THE AMOUNT tony earned. chris earned $60 more than tony. total earned $900

OpenStudy (secret-ninja):

What is your question?

OpenStudy (anonymous):

how much did each person get

OpenStudy (secret-ninja):

Okay. Do you know how to start it off?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

noooooooooooooo

OpenStudy (secret-ninja):

o.o Okay, sheesh. 900 = t (tony) + 2t-40 (jennifer) + t+60 (chris) simplify 900 = 4t+20 880 = 4t simplify 220 = t Tony = $220 Chris = 60+220 = $280 Jennifer = 2t-40 = $400 Check work: 400+280+220= $900

OpenStudy (secret-ninja):

\(\bbox[8pt, pink, border:5pt solid green]{\Large Welcome~to~Openstudy!~:)}\) If this helped, please click the \(\large\sf\color{blue}{Best~Response}\) button and then click the \(\large\sf\color{blue}{Close}\) button to close your question so that you can post a new one later. Go here for all of the basics: http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 Go Here to learn about the rules of Os: http://openstudy.com/code-of-conduct \(\Huge\cal\color{darkgreen}{Have~a~great~day!}\) @krott12589

OpenStudy (anonymous):

thanks you so much